OpenStudy (anonymous):

What is the y coordinate of the y-intercept of the line that passes through the points (-2, -2) and (2, 4)?

OpenStudy (anonymous):

find the slope 4-(-2)=6 2-(-2)=4 6/4=3/2 y=3/2(x-2)+4 y=3/2x+1 y intercept is (0,1)